OpenStudy (anonymous):

Obtain the Laurent series of: \[ f(z) = { z^4 \over 1 + z}, \text{ about z=0 and |z| > 1} \]

OpenStudy (anonymous):

Hint : \[f(z) = { z^4 \over 1 + z}=\frac{z^3}{1+\frac{1}{z}}\]\[|\frac{1}{z}|<1\]

OpenStudy (anonymous):

thanks man!!

OpenStudy (anonymous):

yw :)